OpenStudy (anonymous):

Let f(x)= 16x^5+13x^3+1/√(64x^10+10x^8+2x^4) Find the limit as x→-∞ and x→∞ I know the answers are -2 and 2 now. But when I graph the expression, why does the calculator show the limits are -∞ and ∞?

OpenStudy (experimentx):

divide both sides by +inf ... the limit will be 2 for +inf

OpenStudy (experimentx):

should be -2 for -inf .. i guess .. but how to show it???

OpenStudy (anonymous):

+2 for \( +\infty\)

OpenStudy (anonymous):

-2 for \(-\infty \)

OpenStudy (anonymous):

i havent done anything like this for ages but i think the following approach is ok: multiply by \[\frac{x^{-5}}{x^{-5}}\] to get \[f(x) = \frac{16 + 13x^{-4} + x^{-5}}{\sqrt{64 + 10x^{-2} + 2x^{-6}}}\] now let \[x \rightarrow \infty\] f(x) tends to 2

OpenStudy (anonymous):

is that ok guys?

OpenStudy (experimentx):

yep ..

OpenStudy (anonymous):

You guys are right. How come when I graph the equation, f(x) increases toward infinity?

OpenStudy (anonymous):

@eigenschmeigen: That's amazing! :D

OpenStudy (anonymous):

what about negative, im not sure it works there

OpenStudy (experimentx):

give a try,,, factor our x from denominator.

OpenStudy (anonymous):

How come the graph doesn't back up -2 and 2?

OpenStudy (anonymous):

i have made an error, the numerator is wrong , should be 13 x^(-2)

OpenStudy (experimentx):

the dominant term is x power 5, which means it will be negative as x is negative

OpenStudy (anonymous):

why does my altered function not account for that?

OpenStudy (anonymous):

@eigenschmeigen: It doesn't matter, it's goint to be zero anyways :)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Read the second part of the question now

OpenStudy (anonymous):

still annoys me though. when x is negative after multiplying by (x^-5)/(x^-5) whats happening there?

OpenStudy (experimentx):

it doesn't come that way ... wanna keep trying??

OpenStudy (anonymous):

None of you can answer my real question...

OpenStudy (anonymous):

hang on ill wolfram it

OpenStudy (anonymous):

that will tell us whether its your calculator :)

OpenStudy (anonymous):

did you type it correctly?

OpenStudy (experimentx):

it's minus two .. definitely.

OpenStudy (anonymous):

of-course not.

OpenStudy (anonymous):

Ah, I needed parenthesis for the numerator too. Thank you.

OpenStudy (anonymous):

@FoolForMath , that last remark was not needed.

OpenStudy (anonymous):

i think the key is not to multiply by (x^-5)/(x^-5) but by (x^-4)/(x^-4)

OpenStudy (anonymous):

that leaves x in the numerator

OpenStudy (anonymous):

and the denominator becoomes 8|x|

OpenStudy (anonymous):

which solves our negative problem

OpenStudy (experimentx):

@eigenschmeigen no ... i don't think so.

OpenStudy (anonymous):

@eigenschmeigen , no you factor out x^5 from the numberator and denominator.. you get 16+0+0 / √64+0+0 16/8 and 16/-8 are the answers.. there is no negative problem.

OpenStudy (anonymous):

how do we get the negative limit then?

OpenStudy (anonymous):

aahhh

OpenStudy (experimentx):

take out x^5 out of square root

OpenStudy (anonymous):

yes i see, from the top you factor out x^5 and |x^5| from the bottom

OpenStudy (anonymous):

ok

OpenStudy (experimentx):

yeah thats the trick ... tp add up ... it is safe to assume that the root taken out will be positive (since it is just out of square root) .... and sine function seems to be continuous.

OpenStudy (anonymous):

yeah that was silly of me xD